### v2

# PAD: Phase-Amplitude Decoupling Fusion for Multi-Modal Land Cover Classification

Abstract
The fusion of Synthetic Aperture Radar (SAR) and RGB imagery for land cover classification remains challenging due to modality heterogeneity and underutilized spectral complementarity. Existing methods often fail to decouple shared structural features from modality-complementary radiometric attributes, causing feature conflicts and information loss. To address this, we propose Phase-Amplitude Decoupling (PAD), a frequency-aware framework that separates phase (modality-shared) and amplitude (modality-complementary) components in the Fourier domain, thus reinforcing shared structures while preserving complementary characteristics to improve fusion quality. Unlike prior approaches that overlook the distinct physical properties encoded in frequency spectra, PAD is the first to introduce explicit amplitude-phase decoupling for multi-modal fusion. Specifically, PAD comprises two key components: 1) Phase Spectrum Correction (PSC), which aligns cross-modal phase features via convolution-guided scaling to enhance geometric consistency; and 2) Amplitude Spectrum Fusion (ASF), which dynamically integrates high-frequency and low-frequency patterns using frequency-adaptive multilayer perceptrons, leveraging SAR’s morphological sensitivity and RGB’s spectral richness. Extensive experiments on WHU-OPT-SAR and DDHR-SK datasets demonstrate state-of-the-art performance. Our work establishes a new paradigm for physics-aware multi-modal fusion in remote sensing. The code will be available at https://github.com/RanFeng2/PAD.
Index Terms:
Land Cover Classification, RGB-SAR multi-modality, multi-modal segmentation, Synthetic Aperture Radar, Remote Sensing
I Introduction
Land cover classification (LCC) is a fundamental Earth observation task that categorizes surface materials, such as farmlands, forests, water bodies, and urban areas, using remote sensing imagery. This capability is essential for environmental monitoring, resource management, and urban development [1]. Recently, fusing Synthetic Aperture Radar (SAR) and RGB imagery has emerged as a promising approach to enhance LCC performance: SAR provides robust structural information resilient to atmospheric and illumination variations, while RGB imagery offers rich spectral details that mitigate SAR artifacts such as speckle noise [2, 3].
Figure 1: Motivation and SAR-RGB Spatial-Frequency Difference Analysis. (a) SAR and RGB imagery exhibit modality gaps due to distinct sensing mechanisms (backscattering vs. reflectance). (b) Fusion feature space comprises heterogeneous, shared, and complementary components; effective fusion suppresses heterogeneous features while enhancing shared and complementary patterns. (c)-(g) Frequency-domain analysis on WHU-OPT-SAR (see Section III-A): (c) spatial discrepancies exhibit anisotropic distributions; (d) relative amplitude differences concentrate in high-frequency regions; (e) phase discrepancies are minimal and cluster in low-frequency bands; (f)-(g) sampling-rate variations significantly affect spatial discrepancy distributions.
Despite this complementarity, RGB-SAR fusion faces a key challenge: inherent imaging heterogeneity between microwave backscattering (SAR) and optical reflectance (RGB) leads to significant spectral misalignment and hinders effective cross-modal feature integration (see Fig. 1(a)). In SAR images, objects often appear as discrete backscatter responses, lacking fundamental visual cues, such as geometric contours, textures, and consistent shapes, that are readily available in RGB imagery [4]. Compared to natural-image fusion, integrating SAR and RGB data presents pronounced radiometric and geometric disparities, limiting existing frameworks [5, 6, 7].
Deep learning, particularly convolutional neural networks (CNNs), has achieved notable success in LCC by extracting powerful features [8, 9]. To mitigate modality heterogeneity, most approaches use dual-branch architectures to extract multi-scale features independently, then fuse them via element-wise summation, mean aggregation, or channel-wise concatenation before a shared decoder or segmentation head [10]. However, accurate modeling of complementary features is still impeded by SAR-RGB heterogeneity. Inspired by contrastive learning [11, 12, 13], recent works align modality-shared features to guide complementary feature learning and alleviate inter-modal conflicts (see Fig. 1(b)), yet spatial-domain modeling remains inefficient.
To better capture RGB-SAR complementarity, some methods explore frequency-domain fusion [14, 15, 16]. For example, Li et al. [17] treat phase as modality-shared features but discard amplitude information essential for terrain delineation and do not systematically validate phase sharing.
We conduct comprehensive spectral analysis (see Fig. 1(c)-(g)) to reveal three critical frequency-domain priors: 1) SAR-RGB spatial-domain differences are anisotropic and susceptible to pixel-level mismatch (see Fig. 1(c)); 2) average phase spectra differ minimally and mainly in low frequencies, indicating phase as shared information (see Fig. 1(e)); and 3) amplitude mismatches exhibit distinct high/low-frequency patterns, most pronounced in the high-frequency band, highlighting the need for frequency-aware modeling (see Fig. 1(d)).
Motivated by these insights, we propose Phase-Amplitude Decoupling (PAD), which disentangles phase components capturing modality-invariant structures from amplitude components encoding modality-specific details. By explicitly aligning shared phase representations, PAD enhances complementary amplitude feature extraction and reduces cross-modal conflicts. PAD comprises two modules: Phase Spectrum Correction (PSC), which aligns and reinforces structural details via convolution-guided scaling; and Amplitude Spectrum Fusion (ASF), which dynamically integrates high-frequency and low-frequency patterns using frequency-adaptive MLPs.
Our contributions are fourfold:
  • We systematically analyze registration datasets to establish frequency distributions of SAR-RGB common (phase) and unique (amplitude) patterns, offering interpretable physical priors for multi-source fusion.
  • We introduce frequency-domain amplitude-phase decoupling for multi-modal fusion, orthogonally separating phase (modality-shared) and amplitude (modality-specific) features to resolve heterogeneity-induced conflicts.
  • PSC employs convolution-guided multiplicative scaling to align and reinforce critical structural details for robust fusion.
  • ASF uses frequency-domain MLPs to integrate complementary high-frequency (local details) and low-frequency (global structure) components, leveraging SAR’s morphological sensitivity and RGB’s spectral richness for improved segmentation.
Figure 2: Flowchart of the PAD Framework. Registered RGB and SAR images, $I_{\mathrm{RGB}}$ and $I_{\mathrm{SAR}}$, are processed by asymmetric backbones to extract features $x_{\mathrm{RGB}}$ and $x_{\mathrm{SAR}}$ at each stage. These features are passed through $n$ PAD fusion modules, each comprising: (a) the SCF module, (c) the ASF module, and (d) the PSC module. (b) FD and (e) FR denote the spatial-to-frequency and frequency-to-spatial transformations, respectively. Each stage’s fused features are concatenated channel-wise (outer circle $C$) and fed into a shared decoder. Here, $\mathcal{A}$ and $\mathcal{P}$ represent amplitude and phase spectra.
II Related Work
II-A Multi-modal Fusion for Land Cover Classification
Single-modality approaches [18, 19, 20, 21, 22, 23, 24, 25] perform well on homogeneous data but cannot exploit cross-modal synergies in complex environments. SAR-RGB fusion methods generally follow two paradigms: cross-modal transfer learning and hierarchical feature integration.
II-A1 Cross-Modal Transfer Learning
Transfer learning [26] adapts knowledge from one modality to another, but the pronounced differences between SAR and RGB data often degrade learned representations, limiting multi-modal applicability.
II-A2 Hierarchical Feature Fusion
Feature fusion has evolved through three stages: a) Basic Fusion Operators: Early methods used concatenation or weighted summation [27], but these are prone to modality dominance and redundancy. b) Attention-Guided Fusion: Spatial-channel attention mechanisms address these issues [28, 29]. For example, MCANet’s cross-modal attention (MCAM) models inter-modality dependencies; MRFS [30] introduces progressive cycle attention for infrared-visible fusion; and CMX [31] employs bidirectional channel-spatial attention. However, these paradigms incur high computational costs. c) Complementarity-Aware Fusion: Recent approaches explicitly disentangle shared and modality-specific features. ASANet [32] uses semantic focusing modules to emphasize complementary cues; CDDFuse [12] applies contrastive learning to separate private and shared representations; and CEN [33] leverages dynamic channel exchange with independent batch normalization. Despite these advances, spatial-domain methods still struggle with geometric misalignment and inefficiency.
II-B Frequency Representation Learning in Vision Tasks
Frequency representation learning has emerged as a powerful tool across various vision tasks due to its capacity for global modeling and structural decomposition. In this subsection, we review three relevant areas:
II-B1 Spectral Decomposition for Structure-Style Separation
Fourier-based decomposition isolates phase (structural semantics) and amplitude (appearance) for domain generalization and style transfer [34, 35]. Early methods used feature statistics for style normalization [36, 37], while recent works employ Fourier transforms for clearer separation, enabling content-preserving augmentation by swapping amplitudes across domains [38, 39].
II-B2 Frequency-Driven Global Modeling
Fourier transforms provide a global receptive field useful for efficient attention and context modeling. Fast Fourier Convolution (FFC) [40] and FFTformer [41] leverage spectral processing to replace or accelerate spatial attention. Spectral token mixing has also been effective in lightweight transformers [42, 43].
II-B3 Frequency-Based Inductive Biases in Dense Prediction
Spectral priors enhance dense prediction tasks. DFF [44] uses frequency-domain dynamic filtering to amplify invariant low-frequency features, improving cross-domain generalization. FreqFusion [45] designs learnable high- or low-pass filters for semantic consistency and boundary refinement. FCENet [14] employs frequency-selective modules for cross-modal spectral correlations in RGB-NIR denoising.
In summary, frequency-based representations offer a flexible, physically interpretable design space for vision tasks; yet, few frameworks target task-specific fusion in remote sensing. We address this gap in the next section.
II-C Frequency-Domain Fusion Methods
Applying frequency representations to multi-modal fusion is recent. Shi et al. [15] introduce Frefusion, using spectral attention to merge modalities in the frequency domain. Frequency spectra naturally decompose signals into phase (modality-invariant structures) and amplitude (modality-specific details), offering a principled fusion framework.
However, existing methods often transplant spatial operators, MLPs [16], convolutions [14], or attention [15], into the spectral domain, modulating magnitude while neglecting the distinct roles of phase and amplitude. Consequently, shared information may be preserved, but complementary cues are suppressed.
In contrast, our PAD framework explicitly decouples phase and amplitude spectra. Through systematic spectral analysis (see Section III-A), we show that SAR and RGB share strong phase consistency while amplitude captures complementary patterns. PAD introduces (1) a Phase Spectrum Correction module for robust structural alignment and (2) an Amplitude Spectrum Fusion module with frequency-adaptive refinement. This decomposition-correction-fusion pipeline distinguishes PAD from previous FFT-based approaches by providing a physically interpretable strategy for multi-modal fusion.
The two most related works, FeINFN [16] and SFINet++ [46], also leverage amplitude and phase in the frequency domain for high-frequency detail enhancement and global context modeling. However, they lack a physically grounded interpretation of amplitude-phase roles and do not explicitly decouple shared versus modality-specific information. In contrast, PAD is the first framework to formulate and exploit amplitude-phase decomposition for principled alignment and complementary integration across modalities.
III Proposed Method
To address the fundamental conflict between modality-invariant geometry and modality-specific appearance in SAR-RGB fusion, we introduce Phase-Amplitude Decoupling (PAD), a frequency-aware multi-modal fusion framework. The overall architecture, shown in Fig. 2, comprises three modules: the Spatial Complementary Fusion (SCF) module, which preserves fine-grained spatial priors; the Phase Spectrum Correction (PSC) module, which aligns and reinforces shared structural details; and the Amplitude Spectrum Fusion (ASF) module, which adaptively integrates complementary high- and low-frequency amplitude cues. Unlike previous methods that apply spectral operations uniformly, PAD explicitly decouples and recombines phase and amplitude with dedicated modules to enable physics-consistent fusion.
Input : SAR/RGB image pairs $\{I_{\mathrm{SAR}}^{(i)},I_{\mathrm{RGB}}^{(i)}\}_{i=1}^{N}$
Output : RSD, RAD, APPD metrics for each image pair
1
1: Preprocessing
//Convert RGB image to grayscale
2 $I_{\mathrm{RGB}}\leftarrow 0.299R+0.587G+0.114B$
//Retain SAR image as is
3 $I_{\mathrm{SAR}}\leftarrow I_{\mathrm{SAR}}$
4
2: Spectral Decomposition
5 for each pair $(i)$ do
       //Frequency domain conversion
6       Compute $\mathcal{A}_{\mathrm{SAR}}^{(i)}$ and $\mathcal{P}_{\mathrm{SAR}}^{(i)}$ via Eqs. 11, 12 and 13
7       Compute $\mathcal{A}_{\mathrm{RGB}}^{(i)}$ and $\mathcal{P}_{\mathrm{RGB}}^{(i)}$ similarly
8
      //Spatial difference computation
9       Compute $\Delta I^{(i)}\leftarrow\mathrm{RSD}^{(i)}$ via Eq. 1
       //Amplitude difference computation
10       Compute $\Delta\mathcal{A}^{(i)}\leftarrow\mathrm{RAD}^{(i)}$ via Eq. 2
       //Phase difference computation
11       Compute $\Delta\mathcal{P}^{(i)}\leftarrow\mathrm{APPD}^{(i)}$ via Eq. 3
12
13 end for
14
3: Dataset-level Aggregation
//Mean spatial, amplitude, phase difference computation
15 $\mathrm{RSD}\leftarrow\mathtt{Mean}(\{\Delta I^{(i)}\}_{i=1}^{N})$
16 $\mathrm{RAD}\leftarrow\mathtt{Mean}(\{\Delta\mathcal{A}^{(i)}\}_{i=1}^{N})$
17 $\mathrm{APPD}\leftarrow\mathtt{Mean}(\{\Delta\mathcal{P}^{(i)}\}_{i=1}^{N})$
Algorithm 1 Spectral Difference Analysis.
III-A Frequency-domain Priors for Amplitude-Phase Decoupling
To validate the physical significance of phase-amplitude decoupling and reveal the shared and complementary relations between RGB and SAR, we conducted a systematic spectral analysis on WHU-OPT-SAR dataset, the largest RGB-SAR segmentation benchmark.
Firstly, we visualized the averaged difference spectra (see Fig. 1(c)-(e)) using Algorithm 1 to compare modality differences in spatial, amplitude, and phase domains. This analysis shows that frequency-domain representations capture cross-modal sharing and complementarity more effectively than spatial representations.
The difference spectra are defined as:
$\mathrm{RSD}$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{\left|I_{\mathrm{SAR}}^{(i)}-I_{\mathrm{RGB}}^{(i)}\right|}{I_{\mathrm{RGB}}^{(i)}+\epsilon},$ (1)
$\mathrm{RAD}$ $=\frac{1}{N}\sum_{i=1}^{N}\frac{\left|\mathcal{A}_{\mathrm{SAR}}^{(i)}-\mathcal{A}_{\mathrm{RGB}}^{(i)}\right|}{\mathcal{A}_{\mathrm{RGB}}^{(i)}+\epsilon},$ (2)
$\mathrm{APPD}$ $=\frac{1}{N}\sum_{i=1}^{N}\left|\arg\exp j(\mathcal{P}_{\mathrm{SAR}}^{(i)}-\mathcal{P}_{\mathrm{RGB}}^{(i)})\right|,$ (3)
where $I$, $\mathcal{A}$, and $\mathcal{P}$ denote pixel values, amplitude spectra, and phase spectra for SAR and RGB images; $\arg(\cdot)$ computes the complex angle; and $\epsilon$ prevents division by zero. By leveraging the periodicity of complex exponentials, APPD maps phase differences into $[-\pi,\pi)$, correcting wrap-around artifacts (e.g., $2\pi\to 0$).
Figure 1(c)-(e) reveals three priors: 1) Spatial differences (RSD) exhibit anisotropic distributions. 2) Amplitude differences (RAD) concentrate in high-frequency regions. 3) Phase differences (APPD) remain minimal, with slight low-frequency clustering. These frequency-domain metrics demonstrate superior explainability over spatial RSD, confirming that spectral decomposition better captures cross-modal specificity.
Physically, low-frequency phase disparities stem from geometric distortions (e.g., side-looking SAR vs. orthographic RGB), while high-frequency variations arise from SAR speckle noise. Amplitude disparities reflect modality-specific sensitivity to object scales: low frequencies encode large structures (farmlands, forests) and high frequencies capture fine details (roads).
These analyses yield two key insights that drive our architecture design: 1) Phase Sharing: Consistently low APPD values indicate strong phase consistency. The PSC module thus performs pixel-level alignment while preserving this shared information. 2) Amplitude Complementarity: RAD’s frequency-dependent divergence motivates ASF’s dynamic frequency-radius adjustment to handle amplitude variations.
Secondly, we conducted sensitivity experiments on sampling rates (see Fig. 1(f)-(g)) to evaluate the stability of our spectral metrics. By downsampling input images at various ratios and computing the mean and variance of RSD, RAD, and APPD across the dataset, we observed that RAD and APPD remain nearly constant, whereas RSD exhibits larger fluctuations. This invariance of RAD and APPD to sampling artifacts confirms that frequency-domain complementarity arises from intrinsic modality characteristics rather than from individual changes.
TABLE I: Statistical Analysis of the Average Periodic Phase Difference Spectrum.
Metric Low Frequency High Frequency Whole Image
Shapiro-Wilk Statistic 1.0044 1.0050 1.0131
Shapiro-Wilk p-value 1.0000 1.0000 1.0000
Skewness -0.0521 0.0020 -0.0265
Kurtosis 0.0750 -0.0117 0.0322
Figure 3: Statistical Analysis of APPD. (a) APPD spectrum. (b)-(d) Frequency histograms with Gaussian fits for low-frequency (LF; radius = 0.5 $\times$ half-diagonal), high-frequency (HF), and all spectra (ALL). (e)-(f) Mean-variance trajectories across sampling rates. All spectra exhibit near-Gaussian invariance, with LF and ALL showing slight deviations due to structural noise in the low-frequency range, indicating that phase, particularly HF, can serve as a robust shared feature.
Thirdly, to assess the statistical consistency of cross-modal phase spectra, we performed normality tests (Shapiro-Wilk) and calculated skewness and kurtosis for APPD values at each frequency band (see Table I). Frequency histograms and mean-variance trajectories (see Fig. 3) demonstrate that APPD distributions closely follow a Gaussian profile, especially in high-frequency regions. Low-frequency components show minor deviations, likely due to structural noise, but overall the near-Gaussian behavior validates phase spectra as a reliable shared feature. In a word, phase is so similar between RGB and SAR that it is almost unaffected by individual differences.
Figure 4: Visualization of Spatial and Frequency Spectrum Differences under Missing RGB Regions. (a)-(b) RGB images with missing regions; (c) variation in RSD, RAD, and APPD, showing RSD’s greater sensitivity to missing data; (d)-(f) difference spectra including missing regions; (g)-(i) difference spectra excluding missing regions. These results demonstrate that frequency-domain metrics (RAD, APPD) are more robust to missing data than spatial-domain RSD.
Figure 5: Visualization of Spatial and Frequency Spectrum Differences under Cloud Obscuration. (a)-(c) Example observations with cloud masks derived from RGB, SAR, and NIR channels in WHU-OPT-SAR [47]. (d) Comparison of RSD, RAD, and APPD stability between cloud-covered and cloud-free areas; arrows denote the value of decline.
Finally, we evaluated phase invariance under two challenging scenarios: 1) partial RGB data removal (see Fig. 4) and 2) cloud-obscured observations (see Fig. 5). In the former, RSD increased dramatically when RGB pixels were missing (compare Fig. 4(d) and (g)), while APPD remained stable (compare Fig. 4(f) and (i)). In the latter, APPD variance ($\Delta\approx 10^{-0.037}\%$) was negligible compared to that of RSD ($\Delta\approx 10^{5.81}\%$) and RAD ($\Delta\approx 10^{0.58}\%$) as shown in Fig. 5(d). These results confirm that phase information, particularly at high frequencies, retains robust consistency across modalities, underscoring its role as a shared feature to mitigate inter-modal conflicts and enhance complementary fusion.
III-B Overall Framework
As illustrated in Fig. 2, the proposed PAD framework integrates SAR and RGB modalities via a dual-branch encoder and a shared decoder, with its primary innovation in cross-modal feature fusion through phase-amplitude decoupling. The encoder employs asymmetric convolutional backbones $\mathcal{B}_{\mathrm{SAR}}$ and $\mathcal{B}_{\mathrm{RGB}}$, each with independent parameters to preserve modality-specific attributes. Specifically, $\mathcal{B}_{\mathrm{SAR}}$ extracts microwave scattering features from SAR images, while $\mathcal{B}_{\mathrm{RGB}}$ captures optical reflectance patterns from RGB images. Both backbones produce hierarchical feature maps at four stages with resolutions $\{1/4,1/8,1/16,1/32\}$ for $i\in\{0,1,2,3\}$:
$x_{RGB}^{i}=\mathcal{B}_{\mathrm{RGB}}^{(i)}\left(I_{\mathrm{RGB}}^{i}\right),\quad x_{SAR}^{i}=\mathcal{B}_{\mathrm{SAR}}^{(i)}\left(I_{\mathrm{SAR}}^{i}\right).$ (4)
Although frequency-domain fusion is more robust to modality differences, it can discard rich local interactions in the spatial domain. Therefore, we first introduce a lightweight SCF module to retain fine-grained spatial priors. The SCF module (see Eq. 5) processes the bimodal features by computing pixel-wise fusion weights via spatial attention gating. This operation preliminarily integrates complementary features in the spatial domain and reduces feature dimensionality, constraining the solution space for subsequent frequency-domain operations:
$X^{i}=\mathrm{SCF}\left(x_{RGB}^{i},x_{SAR}^{i}\right),$ (5a)
$\mathcal{S}=\mathrm{Sigmoid}\left(\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(x_{1},x_{2}\right)\right)\right),$ (5b)
$x_{1}^{\prime}=\mathcal{S}\odot x_{1},\quad x_{2}^{\prime}=\left(1-\mathcal{S}\right)\odot x_{2},$ (5c)
$x=\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(x_{1}^{\prime},x_{2}^{\prime}\right)\right).$ (5d)
The fused feature $X^{i}$ is then transformed via FD into separate amplitude $\mathcal{A}^{i}$ and phase $\mathcal{P}^{i}$ spectra:
$\mathcal{A}^{i},\mathcal{P}^{i}=\mathrm{FD}\left(X^{i}\right).$ (6)
The amplitude spectrum $\mathcal{A}^{i}$ is enhanced by the ASF module, which uses a trainable frequency radius $r$ to distinguish and reinforce high-frequency details while preserving low-frequency structures:
$\mathcal{A^{\prime}}^{i}=\mathrm{ASF}\left(\mathcal{A}^{i};r\right),$ (7)
where the PSC module applies multiplicative scaling to correct the phase:
$\mathcal{P^{\prime}}^{i}=\mathrm{PSC}\left(\mathcal{P}^{i}\right).$ (8)
Finally, FR reconstructs the fused features:
$\bm{F^{\prime}}^{i}=\mathrm{FR}\left(\mathcal{A^{\prime}}^{i},\mathcal{P^{\prime}}^{i}\right).$ (9)
The shared decoder $\mathcal{D}$ then aggregates the multi-stage fused features to generate high-precision land-cover segmentation maps:
$\bm{\hat{Y}}=\mathcal{D}\left(\bm{F^{\prime}}^{0},\bm{F^{\prime}}^{1},\bm{F^{\prime}}^{2},\bm{F^{\prime}}^{3}\right).$ (10)
III-C Frequency Decoupling & Frequency Recoupling
To achieve physically disentangled representations in the spectral domain, we propose a bidirectional transformation framework comprising FD and its inverse, FR. First, the input image $X\in\mathbb{R}^{H\times W}$ is transformed from the spatial domain to the frequency domain using a 2D real-input fast Fourier transform (rFFT) [48], yielding the spectrum $\mathcal{F}\in\mathbb{C}^{H\times(\lfloor W/2\rfloor+1)}$:
$\mathcal{F}\left(u,v\right)=\sum_{y=0}^{H-1}\sum_{x=0}^{W-1}X\left(y,x\right)\exp\left(-i2\pi\left(\frac{ux}{W}+\frac{vy}{H}\right)\right),$ (11)
where $(u,v)$ are frequency-domain coordinates. Exploiting the conjugate symmetry of real-valued signals, we retain only the non-redundant region ($0\leq u\leq\lfloor W/2\rfloor$, $0\leq v<H$). Next, an FFT shift relocates low-frequency components to the spectrum center:
$\mathcal{F}_{\mathrm{shift}}\left(u,v\right)=\mathcal{F}\left(u\oplus\frac{W}{2},v\oplus\frac{H}{2}\right),$ (12)
where $\oplus$ denotes modulo addition. Finally, we derive the amplitude and phase spectra as
$\mathcal{A}=|\mathcal{F}|,\quad\mathcal{P}=\angle\mathcal{F},$ (13)
where $\mathcal{A},\mathcal{P}\in\mathbb{R}^{H\times(\lfloor W/2\rfloor+1)}$ capturing radiometric and structural information, respectively. This decoupling allows the model to focus on modality-specific representations: SAR phase carries scattering patterns sensitive to physical surface characteristics, whereas RGB amplitude captures fine-grained textures and semantic content. By separating them, the network avoids early-stage interference between heterogeneous modalities, enhancing feature discriminability before integration.
III-D Phase Spectrum Correction Module
While previous works have assumed perfect phase alignment [17], our spectral analysis (see Fig. 1(e)) reveals that low-frequency geometric inconsistencies persist, necessitating a learnable, spatially aware correction mechanism. Inspired by channel attention, we propose a lightweight Phase Spectrum Correction (PSC) module tailored for frequency-domain phase alignment.
As illustrated in Fig. 1(e), minor misalignments in the low-frequency phase spectra between SAR and RGB, attributable to geometric distortions (e.g., side-looking SAR vs. RGB orthographic projection), can compromise spatial consistency and boundary precision. Traditional rigid registration methods are inadequate for such non-rigid distortions. Thus, the PSC module performs adaptive geometric correction using learnable multiplicative masks.
Unlike the global channel scaling in traditional attention mechanisms, the proposed PSC estimates a spatially varying modulation mask via a stack of $d$ consecutive $1\times 1$ convolutions, enabling localized, fine-grained phase adjustment. Given an input phase-spectrum tensor $P\in\mathbb{R}^{C\times H\times(\lfloor W/2\rfloor+1)}$, we first compress the channel dimension via $d$ $1\times 1$ convolution (reducing $C$ to $C/r$, $r\in\{4,8\}$) followed by leaky linear rectification function (Leaky ReLU) [49]. A second $1\times 1$ convolution produces a single-channel Phase Correction Mask (PCM), constrained to $(0,1)$ by a Sigmoid:
$\mathrm{PCM}=\sigma\left(\mathrm{Conv}_{d}^{1\times 1}\left(\delta\left(\cdots\delta\left(\mathrm{Conv}_{1}^{1\times 1}\left(P\right)\right)\cdots\right)\right)\right),$ (14)
where $d$ denotes the number of stacked convolutions and $\delta(\cdot)$ is Leaky ReLU. Pixel-wise multiplicative scaling is then applied:
$P^{\prime}=P\odot\left(1+\mathrm{PCM}\right),$ (15)
where $\odot$ is the Hadamard product. This formulation enables localized, fine-grained phase adjustment while preserving phase periodicity and incurs minimal computational overhead.
III-E Amplitude Spectrum Fusion Module
Previous frequency-domain methods often apply fixed high- or low-pass filters or spatially uniform spectral attention, failing to adapt to content-specific frequency distributions. In contrast, our analysis in Fig. 1(d) shows that the spectral discrepancy between SAR and RGB modalities is primarily concentrated in the medium-to-high frequency components of the amplitude spectra. To exploit this prior while maintaining gradient stability, we design the ASF module to employ a trainable frequency radius for dynamic separation and enhancement of complementary frequency bands via three sequential operations.
First, we compute a normalized frequency distance metric for each spectral component $i$:
$\mathrm{Dist}(i)=\frac{\sqrt{\left(X_{i}-X_{\mathrm{center}}\right)^{2}+\left(Y_{i}-Y_{\mathrm{center}}\right)^{2}}}{\sqrt{X_{\mathrm{max}}^{2}+Y_{\mathrm{max}}^{2}}},$ (16)
where quantifies component positions relative to the spectrum center $(X_{\mathrm{center}},Y_{\mathrm{center}})$. A trainable boundary radius $r$, constrained via a Sigmoid activation to yield $R=\sigma(r)\in[0,1]$, dynamically partitions the spectrum into low- and high-frequency regions. This constraint matches the normalized spectral dimensions and prevents $R$ from diverging to unreasonable values during training, which would otherwise lead to unstable partitioning and degraded performance.
To maintain differentiability, we employ a soft partition:
$\mathrm{Mask}_{\mathrm{high-freq}}=\sigma\left(\left(\mathrm{Dist}-R\right)\cdot\tau\right),$ (17)
where temperature $\tau=10$. High-frequency components are enhanced via a position-wise MLP:
$A_{\mathrm{high-freq}}=\mathrm{MLP}\left(A\odot\mathrm{Mask}_{\mathrm{high-freq}}\right),$ (18)
then concatenated with the original amplitude $A$ and fused by a $1\times 1$ convolution and residual connection:
$A^{\prime}=\mathrm{Conv}_{1\times 1}\left(\mathrm{Concat}\left(A,A_{\mathrm{high-freq}}\right)\right),\quad A\leftarrow A+A^{\prime}.$ (19)
This module preserves low-frequency structures while adaptively reinforcing high-frequency details that capture complementary semantic information. The learnable radius $R$ self-adjusts to input characteristics, eliminating manual threshold tuning.
III-F Loss Function
Our optimization framework integrates three objectives to jointly promote semantic segmentation accuracy, multi-scale feature learning, and amplitude-phase coherence.
III-F1 Primary Semantic Supervision
The primary segmentation loss is defined as the standard cross-entropy loss:
$\mathcal{L}_{\mathrm{seg}}=-\sum_{i\in\Omega}\sum_{c=1}^{C}y_{i,c}\log\hat{p}_{i,c},$ (20)
where $\Omega$ denotes the set of foreground pixels (excluding background label 255), $y_{i,c}\in\{0,1\}$ indicates the ground-truth class, and $\hat{p}_{i,c}$ is the predicted probability.
III-F2 Auxiliary Semantic Supervision
An auxiliary segmentation head is added at the second-to-last stage ($t=2$), also using cross-entropy loss:
$\mathcal{L}_{\mathrm{aux}}=-\sum_{i\in\Omega}\sum_{c=1}^{C}y_{i,c}\log\hat{p}^{(a)}_{i,c}.$ (21)
III-F3 Phase-Amplitude Decoupling Constraint
To resolve the coupling between amplitude and phase spectra and enforce consistency between fused and original amplitudes, we introduce an amplitude consistency loss that regularizes $\mathcal{A}^{\prime}_{t}$ to remain faithful to its original decoupled state $\mathcal{A}_{t}$ across feature scales:
$\mathcal{L}_{\mathrm{amp}}=\frac{1}{|\mathcal{T}|}\sum_{t\in\mathcal{T}}\alpha_{t}\left\|\mathcal{A}^{\prime}_{t}-\mathcal{A}_{t}\right\|_{F}^{2},$ (22)
where $\mathcal{T}=\{0,1,2,3\}$ indexes the feature scales, $\alpha_{t}=0.5+0.5\frac{t}{3}$ imposes depth-progressive weighting, and $\|\cdot\|_{F}$ is the Frobenius norm.
III-F4 Composite Objective
The overall loss is
$\mathcal{L}_{\mathrm{total}}=\mathcal{L}_{\mathrm{seg}}+\lambda_{1}\mathcal{L}_{\mathrm{aux}}+\lambda_{2}\mathcal{L}_{\mathrm{amp}},$ (23)
with $\lambda_{1}=0.4$ balancing auxiliary supervision and $\lambda_{2}=0.1$ controlling the amplitude regularization strength. This composite loss promotes both geometric invariance, by stabilizing phase-sensitive features through $\mathcal{L}_{\mathrm{amp}}$, and semantic invariance via hierarchical amplitude refinement.
IV Experimental Results
IV-A Dataset Description
WHU-OPT-SAR [28] is a multi-modal dataset for land use classification in Hubei Province, China. It contains 100 co-registered RGB, NIR, and SAR image triples with corresponding land type labels. RGB images are acquired by GF-1, and SAR images by GF-3; both are resampled to 5 m resolution and sized 5556 $\times$ 3704 pixels. Seven primary land use categories are covered: farmland, urban areas, villages, water bodies, forests, roads, and others. We crop each image into non-overlapping 512 $\times$ 512 patches, yielding 7,000 patches, of which 80% (5,600) are used for training and 20% (1,400) for testing.
DDHR-SK is a subset of the DDHR multi-modal dataset for land cover classification in Pohang City, South Korea [29]. RGB images from GF-2 and SAR images from GF-3 are paired, and RGB images are processed with a fog filter in GIMP to simulate adverse conditions. The dataset comprises 6,173 cloud-affected RGB-SAR pairs, resampled to 1 m resolution and resized to 256 $\times$ 256 pixels. It includes five land cover types: buildings, farmland, greenery, water, and roads, with 3,087 training and 3,086 validation images in the public split.
IV-B Training Setting
All models were trained using the MMSegmentation framework [50] on a single NVIDIA H800 GPU (comparisons used an NVIDIA A100). We used a batch size of 4 for 80 k iterations, validating every 8 k iterations. Optimization was performed with AdamW [51], an initial learning rate of $1\times 10^{-4}$, and a cosine annealing schedule (T_max = total iterations). A fixed seed (42) ensured reproducibility.
We report mean IoU (mIoU), Kappa coefficient, overall accuracy (OA), and F1 score, selecting the iteration with the best mIoU. To mitigate class-imbalance bias, we compute the Kappa coefficient per sample and report its mean:
$\kappa_{i}=\frac{P_{o}^{(i)}-P_{e}^{(i)}}{1-P_{e}^{(i)}},\quad\mathrm{mKappa}=\frac{1}{M}\sum_{i=1}^{M}\kappa_{i},$ (24)
where $P_{o}^{(i)}$ and $P_{e}^{(i)}$ are the observed and expected agreements for sample $i$.
IV-C Experimental Results on Datasets
To assess the effectiveness of PAD, we replicated the original configurations of 17 comparison models, nine state-of-the-art multi-modal segmentation approaches and eight leading unimodal methods, on both WHU-OPT-SAR and DDHR-SK.
TABLE II: Quantitative Results (%) on WHU-OPT-SAR. We report per-category IoU for seven land cover classes (Farmland, City, Village, Water, Forest, Road, Others), along with OA, mKappa, mF1, and mIoU. Bold and underlined values indicate the best and second-best performance, respectively.
Method Venue IoU per Category OA mKappa mF1 mIoU
Farmland City Village Water Forest Road Others
RGB DANet [18] CVPR’19 67.51 55.22 46.71 60.85 83.13 38.89 20.43 83.19 49.74 67.40 53.25
OCRNet [19] ECCV’20 67.68 56.48 47.47 61.22 82.98 37.20 17.97 83.23 49.11 66.95 53.00
CGNet [20] TIP’20 61.33 45.96 37.98 49.33 80.98 1.92 0.19 79.26 37.42 50.54 39.67
SegFormer [21] NeurIPS’21 68.18 56.91 46.82 60.85 83.19 35.24 18.73 83.37 49.46 66.79 52.84
SETR [22] CVPR’21 65.25 55.07 42.97 57.80 82.44 31.91 18.36 81.94 46.19 64.73 50.54
Mask2former [23] CVPR’22 67.09 56.64 43.97 60.09 81.47 36.24 21.03 82.69 47.69 66.65 52.36
PIDNet [24] CVPR’23 66.20 55.20 44.95 58.90 82.20 36.36 16.40 82.25 47.20 65.53 51.46
TransUNet [25] MIA’24 67.62 56.28 46.08 60.66 83.25 38.76 19.32 83.18 49.89 67.20 53.14
SAR DANet [18] CVPR’19 60.42 54.12 33.05 54.45 80.57 24.89 15.11 79.44 35.28 60.16 46.09
OCRNet [19] ECCV’20 59.81 52.48 31.21 54.03 79.57 22.49 13.92 78.94 32.62 58.74 44.79
CGNet [20] TIP’20 56.72 46.81 26.32 46.30 77.53 0.27 0.00 76.39 24.64 47.00 36.28
SegFormer [21] NeurIPS’21 60.17 54.71 32.30 54.07 79.79 21.02 15.13 79.31 33.60 59.24 45.31
SETR [22] CVPR’21 60.35 52.29 31.41 53.56 80.15 18.66 15.18 79.12 34.25 58.33 44.51
Mask2former [23] CVPR’22 60.39 52.86 35.69 53.85 78.36 29.16 17.20 79.26 32.76 61.35 46.79
PIDNet [24] CVPR’23 58.59 52.35 29.65 53.60 79.35 23.06 13.74 78.42 32.15 58.33 44.34
TransUNet [25] MIA’24 61.93 54.85 33.42 56.26 80.63 27.56 14.96 80.14 37.45 61.13 47.09
SAR & RGB MCANet [28] JAG’21 64.21 45.49 42.93 56.20 81.81 25.93 8.47 80.80 45.47 59.94 46.44
CEN [33] TPAMI’22 66.46 56.28 45.46 58.09 82.34 32.51 17.82 82.32 47.47 65.36 51.28
TokenFusion [52] CVPR’22 67.87 57.26 46.96 61.66 83.31 36.34 18.76 83.38 49.87 67.09 53.16
DDHRNet [29] JAG’22 59.12 45.67 30.77 47.96 79.59 0.47 0.03 78.00 27.18 48.36 37.66
CDDFuse [12] CVPR’23 59.88 35.08 25.22 52.31 78.77 17.64 15.00 77.02 38.69 54.29 40.56
CMX [31] TITS’23 69.24 58.20 48.45 64.06 83.83 39.99 21.92 84.18 52.76 69.01 55.10
MRFS [30] CVPR’24 69.39 58.13 49.43 64.90 83.87 41.26 21.72 84.21 54.33 69.38 55.53
FTransUNet [6] TGRS’24 64.87 53.84 43.71 55.63 81.53 0.00 0.00 81.25 43.53 52.97 42.80
ASANet [32] ISPRS’24 69.18 56.95 48.58 64.06 83.79 41.48 21.74 84.15 52.62 69.05 55.11
PAD (Ours) 69.85 58.10 49.70 65.39 84.05 43.24 23.47 84.56 53.93 70.14 56.26
IV-C1 Evaluations on WHU-OPT-SAR
As shown in Table II, PAD achieves new state-of-the-art results: 56.26% mIoU, 84.56% OA, 53.93% mKappa, and 70.14% mF1. This surpasses the best unimodal method (DANet-RGB [18]: 53.25% mIoU, 83.19% OA) and the multi-modal baseline MRFS [30] (55.53% mIoU, 84.21% OA). In fine-grained classes, PAD attains 43.24% IoU on roads and 49.70% on villages, improving over MRFS by 1.98% and 0.27%, respectively.
Figure 6: Visualization Results on WHU-OPT-SAR. Four representative scenarios: (a) building-dense area, (b) specular water bodies, (c) vegetated region, and (d) urban-rural transition. For each scenario, images are shown (left to right): RGB, SAR, pseudo-color ground truth, and prediction maps from six competing methods. We compare PAD against single-modal baselines (DANet-RGB, TransUNet-SAR) and multi-modal approaches (CDDFuse, CMX, MRFS). Regions of interest are highlighted with gray boxes (buildings/water) and circles (vegetation); see Section IV-C for details.
Visual results in Fig. 6 further illustrate PAD’s benefits. In (a), PAD recovers occluded road segments (cyan) in building-dense areas via SAR phase cues, whereas RGB-only models miss them. In (b), PAD distinguishes specular reflections from true water boundaries under cloud interference. In (c)-(d), PAD resolves vegetation spectral ambiguities and maintains consistency across urban-rural transitions.
TABLE III: Quantitative Results (%) on DDHR-SK. We report per-category IoU for seven land cover classes (Buildings, Roads, Greenery, Water, Farmland), along with OA, mKappa, mF1, and mIoU.
Method Venue IoU per Category OA mKappa mF1 mIoU
Buildings Roads Greenery Water Farmland
RGB DANet [18] CVPR’19 94.46 75.11 91.10 99.03 90.59 96.23 73.35 94.57 90.06
OCRNet [19] ECCV’20 90.18 60.49 84.42 97.87 82.84 93.17 58.96 90.26 83.16
CGNet [20] TIP’20 80.19 27.57 72.77 90.47 69.97 86.16 33.90 78.76 68.19
SegFormer [21] NeurIPS’21 92.44 67.11 88.76 98.82 88.18 95.00 66.73 92.71 87.06
SETR [22] CVPR’21 94.17 74.10 91.43 99.18 90.88 96.22 73.79 94.49 89.95
Mask2former [23] CVPR’22 94.16 80.46 91.08 98.68 86.88 96.07 74.05 94.76 90.25
PIDNet [24] CVPR’23 87.77 59.01 80.57 96.89 77.64 91.44 53.43 88.56 80.38
TransUNet [25] MIA’24 90.67 62.68 85.64 98.41 83.71 93.64 61.82 90.95 84.22
SAR DANet [18] CVPR’19 91.20 58.97 85.93 98.13 82.31 93.57 59.35 90.28 83.31
OCRNet [19] ECCV’20 86.38 38.70 78.36 96.18 71.34 89.52 41.07 83.54 74.19
CGNet [20] TIP’20 81.16 19.58 69.45 90.27 60.78 84.63 25.42 74.96 64.25
SegFormer [21] NeurIPS’21 88.81 46.92 81.99 97.57 76.75 91.51 49.55 86.73 78.41
SETR [22] CVPR’21 90.97 59.57 84.71 98.16 80.51 93.27 59.39 89.99 82.79
Mask2former [23] CVPR’22 91.34 58.52 86.40 93.29 80.08 93.20 62.69 89.49 81.92
PIDNet [24] CVPR’23 84.37 36.77 73.28 94.91 66.91 87.55 38.63 81.49 71.25
TransUNet [25] MIA’24 86.56 38.45 77.56 96.53 71.30 89.49 42.21 83.44 74.08
SAR & RGB MCANet [28] JAG’21 85.12 47.88 78.69 96.96 73.18 89.59 46.13 85.55 76.37
CEN [33] TPAMI’22 91.78 58.73 87.70 98.33 85.73 94.26 60.19 90.93 84.46
TokenFusion [52] CVPR’22 93.90 70.67 90.62 98.94 89.21 95.79 69.64 93.70 88.67
DDHRNet [29] JAG’22 79.78 13.85 69.85 93.73 62.17 84.73 24.59 73.75 63.88
CDDFuse [12] CVPR’23 96.59 83.02 94.59 99.11 93.30 97.61 80.65 96.46 93.32
CMX [31] TITS’23 96.43 82.59 94.37 99.38 93.72 97.58 81.08 96.44 93.30
MRFS [30] CVPR’24 96.48 83.68 94.48 99.39 93.84 97.65 82.24 96.60 93.57
FTransUNet [6] TGRS’24 84.69 35.21 76.51 96.40 71.11 88.56 36.55 82.35 72.78
ASANet [32] ISPRS’24 95.04 77.09 92.64 99.24 91.82 96.73 75.77 95.21 91.17
PAD (Ours) 97.09 86.12 95.38 99.47 94.89 98.04 84.38 97.16 94.59
IV-C2 Evaluations on DDHR-SK
As shown in Table III, PAD exhibits superior performance under adverse (cloudy) conditions on DDHR-SK. It achieves 94.59% mIoU and 97.65% OA, outperforming leading unimodal methods (e.g., Mask2former-RGB: 90.25% mIoU, 94.76% mF1; DANet-SAR: 83.31% mIoU, 90.28% mF1) and the multi-modal baseline MRFS (93.57% mIoU, 82.24% mKappa). PAD also attains 82.24% mKappa and 96.60% mF1, with boundary-sensitive categories achieving 99.47% IoU on water and 86.12% IoU on roads, surpassing MRFS by 2.44% and 0.08%, respectively. Qualitative results in Fig. 7 further demonstrate PAD’s ability to delineate water infiltration in irrigated fields and maintain continuous road networks beneath fog.
Figure 7: Visualization Results on DDHR-SK. Four boundary-focused scenarios: (a) coastal demarcation, (b) water-cropland interface, (c) vegetation-building complex, and (d) road network topology. For each scenario, panels show the RGB input, SAR coherence map, pseudo-color ground truth, and prediction maps from six methods. We compare PAD against unimodal baselines (Mask2former-RGB, DANet-SAR) and multi-modal approaches (CDDFuse, CMX, MRFS). Critical boundary regions are highlighted with orange boxes; see Section IV-C for details.
The visual results further demonstrate PAD’s capability for cross-modal boundary refinement. In Fig. 7(a), PAD accurately reconstructs periodically submerged coastlines via SAR coherence analysis, whereas RGB-only models confuse tidal sediments with terrestrial areas. In Fig. 7(b), PAD leverages SAR features to delineate water-infiltration boundaries in irrigated fields, overcoming the over-segmentation common in optical sensors. Moreover, in Fig. 7(c) and Fig. 7(d), PAD achieves continuous road networks under vegetation occlusion and precise material boundaries in vegetation-building complexes, respectively.
TABLE IV: Computational Efficiency Comparison. Model size (#Params), inference time per image, and throughput (FPS) for competitive methods. Lower #Params and inference time, and higher FPS, indicate better efficiency.
Method #Params (M) Inference Time (ms) FPS (frames/s)
CMX [31] 15.00 206.70 1,957.8
MRFS [30] 10.30 695.85 1,342.2
ASANet [32] 7.82 72.91 2,245.2
PAD (Ours) 3.88 5.16 3,228.7
IV-C3 Computational Efficiency Analysis
We evaluate the computational efficiency of PAD in terms of model size (Fusion Modules Only), inference time per image, and frames per second (FPS), as summarized in Table IV. We compare against four high-mIoU methods, CMX [31], MRFS [30], ASANet [32], and CDDFuse [12] (excluded from detailed analysis due to its two-stage design and lower efficiency). All metrics are aggregated across fusion stages.
As shown in Table IV, PAD uses only 3.882 M parameters, achieves the fastest inference time (5.16 ms per image), and delivers the highest FPS (3,228.7). These results demonstrate that PAD not only matches or exceeds accuracy of existing methods but also offers real-time efficiency suitable for operational deployment.
IV-C4 Ablation Study
We performed comprehensive ablation studies on WHU-OPT-SAR to validate the contributions of our core components: the Amplitude Spectrum Fusion (ASF) module, the Phase Spectrum Correction (PSC) module, and Multi-Stage Fusion (MSF).
[…]
ID Components IoU per Category OA mKappa mF1 mIoU
MSF ASF PSC Farmland City Village Water Forest Road Others
[…]
Main Modules
Eight variants (see Table V) were evaluated. The baseline (I), which replaces MSF with early fusion and substitutes PAD (ASF+PSC) with pixel-wise addition, achieves the lowest mIoU (54.98%). Enabling MSF alone (II) yields a +0.39% gain, ASF alone (III) reaches 55.86% mIoU, and PSC alone (IV) provides a smaller improvement. Combining MSF with ASF (V) or PSC (VI) yields similar gains (56.11-56.12% mIoU). Integrating all three modules achieves the best performance (56.26% mIoU), demonstrating their effective synergy.
TABLE VI: Ablation Study on Fusion Modules (%) for WHU-OPT-SAR. We compare alternative amplitude-only, phase-only, and full fusion designs by integrating each into our framework.
Fusion Module IoU per Category OA mKappa mF1 mIoU
Farmland City Village Water Forest Road Others
Amplitude FelNFN [16] 66.93 55.80 45.15 61.27 82.77 35.54 19.17 82.95 47.31 66.46 52.38
FCENet [14] 68.27 55.18 46.79 61.17 83.43 37.09 16.71 83.43 49.82 66.52 52.66
FGSwin [53] 69.87 57.43 49.03 65.04 84.12 43.16 22.26 84.51 53.53 69.70 55.85
Ours 69.91 58.51 49.53 65.52 84.05 42.63 22.58 84.55 54.53 69.93 56.11
Phase FelNFN [16] 60.66 53.25 31.99 54.14 80.64 19.84 15.02 79.42 33.32 58.89 45.08
Channel Attention 69.58 58.66 49.78 64.28 83.96 43.43 22.65 84.39 53.54 69.93 56.05
Ours 69.98 57.32 49.36 65.62 84.12 43.42 23.01 84.57 53.80 69.98 56.12
Whole FelNFN [16] 66.43 51.97 44.36 58.48 82.80 32.33 12.64 82.33 47.35 63.63 49.86
FCENet++ [14] 66.87 53.01 46.41 60.30 82.64 30.21 14.51 82.78 43.83 64.33 50.56
FGSwin [53] 69.88 57.86 49.48 65.26 84.15 42.49 22.36 84.56 53.41 69.76 55.92
Frefusion [15] 69.91 57.99 49.35 64.22 84.00 42.72 22.46 84.41 53.98 69.69 55.81
SFINet++ [46] 68.23 55.65 47.31 62.35 83.10 38.48 19.04 83.50 49.56 67.43 53.45
Ours 69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
Fusion Modules
To further validate ASF and PSC, we replaced each with five existing frequency-domain designs, FelNFN [16], FCENet [14], FGSwin [53], Frefusion [15], and SFINet++ [46], while keeping the rest of our framework unchanged (see Table VI). These methods represent common strategies such as fixed-kernel spectral convolution (FelNFN), frequency-spatial interaction via Hadamard products/matrix multiplication (FCENet), learned spectral filtering (FGSwin), attention-based cross-domain fusion (Frefusion), and amplitude-phase parallel processing (SFINet++). Our ASF outperforms other amplitude-specific methods (56.11% vs. 55.85% mIoU for FGSwin), and PSC consistently exceeds phase baselines (56.12% vs. 56.05% for channel attention and 45.08% for FelNFN). At the full-fusion level, our integrated design still leads (56.26% mIoU), indicating both modules are effective in isolation and synergistic when combined. These results confirm that adaptive frequency modulation (ASF) and spatially guided phase correction (PSC) offer structurally unique and more effective alternatives to existing paradigms.
TABLE VII: Ablation Study on Loss Function Weights (%) for WHU-OPT-SAR. We vary $\lambda_{1}$ with fixed $\lambda_{2}=0.1$ and $\lambda_{2}$ with fixed $\lambda_{1}=0.4$.
Configuration IoU per Category OA mKappa mF1 mIoU
Farmland City Village Water Forest Road Others
Fixed $\lambda_{2}=0.1$
$\lambda_{1}=0.01$ 69.87 57.71 49.43 65.31 84.12 43.35 23.22 84.54 53.60 70.02 56.14
$\lambda_{1}=0.4$ 69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
$\lambda_{1}=1.0$ 69.56 57.30 49.10 65.39 83.84 41.59 20.97 84.37 52.20 69.21 55.39
Fixed $\lambda_{1}=0.4$
$\lambda_{2}=0.01$ 69.83 56.80 48.67 65.63 84.19 43.17 22.82 84.52 53.54 69.76 55.87
$\lambda_{2}=0.1$ 69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
$\lambda_{2}=1.0$ 69.83 57.74 49.37 64.95 83.95 43.45 22.50 84.46 53.87 69.84 55.97
Loss Function Weights
We conducted weight ablations (see Table VII) to assess the trade-off between the amplitude consistency loss $\mathcal{L}_{\mathrm{amp}}$ and segmentation loss $\mathcal{L}_{\mathrm{seg}}$. The best mIoU (56.26%) is achieved with $\lambda_{1}=0.4$ and $\lambda_{2}=0.1$. Increasing $\lambda_{1}$ above 0.4 degrades performance, especially on under-represented classes, while too small $\lambda_{2}$ significantly reduces accuracy, confirming segmentation supervision must remain dominant.
TABLE VIII: Ablation Study on Initial Frequency Radius (%) for WHU-OPT-SAR. We vary the trainable radius parameter in the ASF module.
Radius IoU per Category OA mKappa mF1 mIoU
Farmland City Village Water Forest Road Others
$0.01$ 70.08 57.44 49.31 66.29 84.18 43.08 22.25 84.66 53.92 69.88 56.09
$0.1$ 69.85 58.09 49.71 65.39 84.05 43.24 23.47 84.56 53.96 70.14 56.26
$1$ 69.79 58.19 49.71 64.81 84.15 43.44 23.03 84.52 53.88 70.04 56.16
$10$ 69.71 57.71 49.69 65.63 84.18 43.21 23.56 84.57 53.93 70.12 56.24
$100$ 69.90 57.23 49.57 65.23 84.13 43.34 22.94 84.53 53.88 69.93 56.05
Initial Frequency Radius
We varied the initial frequency radius $r$, which controls the preserved low-frequency proportion in ASF (see Table VIII). The optimal balance is at $r=0.1$ (56.26% mIoU). Values that overly suppress or over-smooth low-frequency content degrade performance, although the method remains robust for $r\in[0.1,10]$.
IV-D Visualization at Distribution Level
To further analyze feature distributions, we extracted multi-scale features from the RGB and SAR encoders and from the fusion modules. After upsampling to the original resolution, we concatenated features channel-wise. To balance computational efficiency with discriminative power, we applied a two-stage dimensionality reduction: a) random spatial sampling, and b) principal component analysis (PCA), retaining 50 components (over 89% variance). The final visualization was generated using t-SNE [54] (perplexity = 30, 1,000 iterations).
Refer to caption Refer to caption Refer to caption
(a) Before PAD Fusion (RGB branch) (b) Before PAD Fusion (SAR branch) (c) After PAD Fusion
Figure 8: t-SNE Visualization of Cross-modal Feature Distribution Evolution. (a) RGB encoder features show spectral continuity but blurred class boundaries. (b) SAR encoder features capture geometric details but are affected by speckle noise. (c) PAD-fused features exhibit enhanced intra-class compactness and improved inter-class separation.
As shown in Fig. 8, PAD-fused features exhibit enhanced intra-class compactness and improved inter-class separability. Quantitatively, mean intra-class standard variance is reduced by 20.39% for roads, 13.16% for water, 6.46% for forest, and 4.43% for farmland, compared to the best unimodal baselines. Moreover, Euclidean distances between critical class pairs increase significantly: City/Village by 75.75% (66.56 $to$ 116.95) and Farmland/Forest by 13.29% (107.64 $to$ 121.89), indicating that PAD captures cross-modal discriminative patterns effectively.
TABLE IX: Statistical Comparison Between Models w/ and w/o ASF.
Metric w/ ASF w/o ASF
Mean Pearson Correlation 0.0364 0.0369
Mean Spearman Correlation 0.0482 0.0488
Paired t-test p-value 0.0016
Wilcoxon test p-value $6.4\times 10^{-7}$
Cohen’s $d$ (effect size) -0.038
IV-E Noise Sensitivity of ASF
To assess whether ASF amplifies high-frequency speckle noise in SAR images, we conducted a correlation analysis using the Equivalent Number of Looks (ENL), a standard measure of speckle. For each image, we computed pixel-wise ENL maps and then measured the Pearson and Spearman correlations between ENL and classification accuracy. As shown in Table IX, the model with ASF exhibits nearly identical correlation coefficients to the baseline, differences in mean correlation are below 0.001, and Cohen’s $d=-0.038$ indicates a negligible effect size. These results confirm that ASF does not increase noise sensitivity and remains robust in low-SNR regions.
Overall, these results confirm that PAD significantly improves classification accuracy and prediction consistency by integrating multi-modal information, suppressing noise, and preserving discriminative spectral and geometric characteristics inherent in the data.
V Conclusion
In this paper, we proposed the Phase-Amplitude Decoupling (PAD) framework for fusing SAR and RGB modalities in land cover classification. Our frequency-domain analysis quantitatively demonstrates that phase components remain consistent across modalities, while amplitude spectra capture modality-specific details, providing the first physical prior for multi-source remote sensing fusion. PAD disentangles and aligns phase features via the Phase Spectrum Correction (PSC) module, and adaptively fuses amplitude components using the Amplitude Spectrum Fusion (ASF) module, which integrates SAR’s high-frequency details with RGB’s low-frequency context. Extensive experiments on two challenging datasets show that PAD outperforms existing methods in both accuracy and robustness.
Despite its demonstrated effectiveness, PAD assumes well-aligned input modalities and relies on hard-label supervision. In future work, we will integrate self-supervised cross-modal registration refinement and soft-label modeling, and investigate transformer-based frequency fusion to further enhance global context awareness.